In Desmos: Write the domain in inequality notion y= 2 square root of x-2+4

1
Expert's answer
2021-03-10T07:09:07-0500

Function is given as:

"y=\sqrt[2]{x^{-2}+4}"


"y=\sqrt[2]{\frac{1}{x^2}+4}"


for real solutions of y the term inside the square root should be as:

"\frac{1}{x^2}+4\ge0"


We can se that x has degree two so for all values of x the above term gives always positive values except x=0, as this makes it infinite value i.e. undefined, so the domain of the given equation is as:

Domain = ("-\infty,+\infty" )-[0]